a man pulls a roller by a force of 20 kgf applied at 60 degree with the ground.if he pulls it a distance of 10 metre in 1 minute .calculate the power dissipated

Grade:11

2 Answers

Trishna mondal
5 years ago
588 kilowat
 
 
riya Dutta
4 years ago
Power=Fcos60*10/time=20*10*1/2*10/60=100/6=16.67Watt
This is because the component of force is Fcos60 along the direction of motion 
Also F is to be expressed in Newton so 20*g or 20*10